A copper wire has a resistance of πŸπŸ“. 𝟎Ω at a temperature of 14℃. After a




current passes through the windings, the resistance rose to 𝟏𝟐. πŸ“Ξ©. To what




temperature was the wire heated?

1
Expert's answer
2022-03-07T12:43:37-0500

We know that

"R=R_0(1+\alpha(T_2-T_1))"


"15=12.5(1+3.93\times10^{-3}(14-T_0))"

"1.2-1=3.93\times10^{-3}(14-T_0)"

"50.89=T_0-14\\T_0=64.89°C"